The diameter of a deoxyribonucleic acid (DNA) molecule is about 20 angstroms. Express the surface area of the molecule in a square meters and square inches. 


1
Expert's answer
2021-06-29T13:15:53-0400

We can treat the deoxyribonucleic acid (DNA) molecule as a sphere. Then, we can find the surface area of the molecule as follows:


"A=4\pi r^2."

Let's first find the surface area of the molecule in square meters:


"A=4\pi\cdot (0.5\cdot20\ \text{\AA}\cdot \dfrac{1.0\cdot10^{-10}\ m}{1\ \text{\AA}} )^2=1.26\cdot10^{-17}\ m^2."

Let's also find the surface area of the molecule in square inches:


"A=4\pi\cdot (0.5\cdot20\ \text{\AA}\cdot \dfrac{3.937\cdot10^{-9}\ in}{1\ \text{\AA}} )^2=1.95\cdot10^{-14}\ in^2."